A jet slows uniformly from 153 km/h to 0 km/h over 42 m. What is the jet's accelertion?
1
Expert's answer
2020-03-11T10:28:02-0400

As per the given question,

Initial velocity of the jet "(u)=153 km/hour=\dfrac{153\times1000}{3600}m/sec=42.5 m/sec"

Final velocity of the jet (v)=0

Distance covered (d)=42m

Let the acceleration of the jet is a,

We know that , "v^2=u^2-2ad"

"\Rightarrow 0=42.5^2-2a\times42"

"\Rightarrow a=\dfrac{42.5\times 42.5}{2\times 42}=21.5 m/sec^2"

"a=21.5m\/sec^2"